# Intrinsic inverse band gap versus polarization relation in ferroelectric materials

### v1

@@ Abstract
Figure 6: Isostructural ferroelectrics (a) AlHO2 and (b) CsAs(HO2)2 displaying an indirect band gap polarization relation under the effect of isotropic strain. For each ferroelectric, we report band gap (eV; purple), Berry phase polarization ($\mu C/cm^{2}$; orange) and Born effective charges polarization ($\mu C/cm^{2}$; green)
@@
Isostructural materials are identified within the Materials Project by imposing structural conditions, specifically, equal space group, total number of atoms, number of species, Wyckoff positions, and coordination number around each atom. These conditions are screened using the pymatgen functions Ong et al. (2013). Our search identifies 21 polar and 25 non polar materials in the Materials Project isostructural to one of the prototypical materials in Table 2. However, only 9 polar and 6 non polar are insulating. We perform full structural relaxations and compute band gaps using dense-grid density of states calculations for these 15 materials. For the case of polar structures, the associated reference non-polar structure is constructed isostructural to the non-polar structure in the ferroelectric database. In this case, we confirm the non polar structure is insulating, therefore allowing a direct calculation of the Berry phase polarization. Similarly, for the case of the non-polar structures, the polar structure is constructed isostructural to the polar structure in the ferroelectric database. In this case, the stability of the polar structure is determined from full structural relaxations.
@@
We confirm 5 cases with a band gap narrowing from non-polar to polar, namely AlHO2 Pmn21, Cd(NO3)2 Pca21, CsAs(HO2)2 Fdd2, YCoO3 P63cm, and YMnO3 P63cm. Note that YMnO3 is a well-known geometric ferroelectric Van Aken et al. (2004). From these, only AlHO2, CsAs(HO2)2 display and inverse band gap polarization relation (see Fig. 6). Table 2 reports their Berry phase polarization, PBE band gaps, as well as energy and volume differences between non-polar and polar phases. CsAs(HO2)2 is a well known ferroelectric Lines and Glass (1977), and the polar structure of AlHO2 is experimentally accessible. Notably, AlHO2 Pmn21 and CsAs(HO2)2 Fdd2 display larger polarizations and smaller PBE band gaps than their isostructures counterparts InHO2 Pmn21 and KP(HO2)2 Fdd2, respectively.
@@
Table 3: Investigation survey that shows the process between the starting materials found in database, calculated materials with decreasing band gap and increasing polarization, monotonous band gap narrowing materials, isostructural candidates, and the final number of materials that exhibit decreasing band gap and increasing polarization
@@
Ferroelectrics Ferroelectric database 200
database Ferroelectrics with $E^{P}_{g}<E^{NP}_{g}$ 20
Inverse $E_{g}$ vs. $P$ relation 9
Materials Polar isostructures 21
Project Non polar isostructures 25
Isostructures with $E^{P}_{g}<E^{NP}_{g}$ 4
Inverse $E_{g}$ vs. $P$ relation 2
Interestingly, the ferroelectrics in Table 2 were identified as inverse band gap versus polarization ferroelectrics simply due to the fact that their prototypical ferroelectric class representative displayed themselves the property. The latter suggest that other inverse relation ferroelectrics may exist within families of ferroelectrics whose representative in the ferroelectric database does not display the property. The latter calls for a systematic search of materials with the property.
In usual ferroelectrics where the polarization arises due to chemical bonding or orbital hybridization, the behavior of the band gap is determined by polarization orientation, symmetry breakings, volume, and type of near edge orbital contribution. Here, ferroelectrics are characterized by low conduction density of states and valence orbital hybridization, therefore band gaps and polarizations are greatly affected by the volume. We emphasize that the band gap reduction from non-polar and polar structures $\Delta E_{g}>0$, is not a sufficient condition to identify inverse relation ferroelectrics. In addition, not all ferroelectrics within a certain class display the inverse relation property, therefore the property is independent of the ferroelectric mechanism. We conclude that in addition to low density of states contribution at the conduction states and negligible orbital hybridization at the valence states, inverse relation ferroelectrics require the right rates of band gap and polarization change under the external parameters, which in turn is determined by details of orbital interaction and distortions in the polar structure.
In summary, we have found 11 examples of ferroelectric materials displaying a novel property, an inverse relation between band gap and polarization. We find 9 examples of such type of ferroelectrics within the ferroelectric database and propose 2 candidates present in the Materials Project. Our results suggest new pathways to engineer small-band gap large-polarization ferroelectrics.
@@ I Acknowledgments
This work was supported by ANID FONDECYT Regular grant number 1220986. Powered@NLHPC: This research was supported by the supercomputing infrastructure of the NLHPC (CCSS210001).

### v2

@@ Abstract
Figure 6: Isostructural ferroelectrics (a) AlHO2, (b) CsAs(HO2)2, (c) GaHO2, (d) Mg3B7BrO13, (e) YHO2 and (f) Zn3B7BrO13 displaying an indirect band gap polarization relation under the effect of isotropic strain. For each ferroelectric, we report band gap (eV; purple), Berry phase polarization ($\mu C/cm^{2}$; orange) and Born effective charges polarization ($\mu C/cm^{2}$; green)
@@
Isostructural materials are identified within the Materials Project by imposing structural conditions, specifically, equal space group, total number of atoms, number of species, Wyckoff positions, and coordination number around each atom. These conditions are screened using the pymatgen functions Ong et al. (2013). Our search identifies 59 polar and 71 non-polar materials in the Materials Project isostructural to one of the prototypical materials in Table 2. However, only 42 polar and 41 non-polar are insulating. From these, 20 polar and 6 non-polar cases contain lanthanides ions, whereas 12 polar and 10 non-polar cases contain magnetic ions with open $d$ orbitals. We disregard these cases since they would require magnetic calculations with multiple spin configurations and computationally intensive calculations that are out of the scope of the work. We perform full structural relaxations and compute band gaps using dense-grid density of states calculations for the remaining 10 polar and 25 non-polar materials. For the case of polar structures, the associated reference non-polar structure is constructed isostructural to the non-polar structure in the SFD. In this case, we confirm the non-polar structure is insulating, therefore allowing a direct calculation of the Berry phase polarization. Similarly, for the case of the non-polar structures, the polar structure is constructed isostructural to the polar structure in the SFD. In this case, the stability of the polar structure is determined from full structural relaxations.
@@
We confirm 7 cases with a band gap narrowing from non-polar to polar, namely XHO2 X = Al, Ga, Y, Cd(NO3)2, CsAs(HO2)2, Mg3B7BrO13 and Zn3B7BrO13. From these, we find that the polar structures of XHO2 X = Al, Ga, Y, CsAs(HO2)2, Mg3B7BrO13 and Zn3B7BrO13 display an inverse band gap polarization relation (see Fig. 6). Table 2 reports their Berry phase polarization, PBE band gaps, as well as energy and volume differences between non-polar and polar phases. CsAs(HO2)2 Lines and Glass (1977) and Zn3B7BrO13 Campa-Molina et al. (2006) are well-known ferroelectrics, and the polar structures of AlHO2 and GaHO2 are experimentally accessible. Notably, CsAs(HO2)2 Fdd2 displays larger polarization and smaller PBE band gap than its isostructural counterpart KP(HO2)2 Fdd2.
@@
Table 3: Number of materials in the ferroelectric database of Smidt et al. and the Materials Project database, materials with band gap narrowing between non-polar and polar structures, isostructural candidates, and the final candidates that exhibit an inverse band gap versus polarization relation.
@@
Ferroelectrics Total database 200
database Ferroelectrics with $E^{P}_{g}<E^{NP}_{g}$ 19
of Smidt et al. Inverse $E_{g}$ vs. $P$ relation 9
Materials Polar isostructures 59
Project Non polar isostructures 71
database Isostructures with $E^{P}_{g}<E^{NP}_{g}$ 7
Inverse $E_{g}$ vs. $P$ relation 6
Interestingly, the ferroelectrics in Table 2 were identified as inverse band gap versus polarization ferroelectrics simply due to the fact that their prototypical ferroelectric class representative displayed the property themselves. The latter suggests that other inverse relation ferroelectrics may exist within families of ferroelectrics whose representative in the SFD does not display the property. The latter calls for a systematic search for materials with the property.
In usual ferroelectrics where the polarization arises due to chemical bonding or orbital hybridization, the behavior of the band gap is determined by polarization orientation, symmetry breakings, volume, and type of near edge orbital contribution. Here, ferroelectrics are characterized by low conduction density of states and valence orbital hybridization; therefore, band gaps and polarizations are greatly affected by the volume. We emphasize that the band gap reduction from non-polar and polar structures $\Delta E_{g}>0$ is not a sufficient condition to identify inverse relation ferroelectrics. In addition, not all ferroelectrics within a certain class display the inverse relation property; therefore the property is independent of the ferroelectric mechanism. We conclude that in addition to low density of states contribution at the conduction states and negligible orbital hybridization at the valence states, inverse relation ferroelectrics require the right rates of band gap and polarization change under the external parameters, which in turn is determined by details of orbital interaction and distortions in the polar structure.
Ferroelectrics with a small band gap and large polarization have potential applications as photovoltaics. Our results suggest that YBiO3 is a promising candidate (see Table 1). Table S6 shows the effective masses computed for the band structures in Fig. 5. The ferroelectric materials in Table I tend to display large (small) effective masses at the VBM (CBM), suggesting relatively poor (good) hole (electron) transport. The study of optical and transport properties of these materials is left for future work.
In summary, we have found 15 examples of ferroelectric materials displaying a novel property, an inverse relation between band gap and polarization. We find 9 examples of such type of ferroelectrics within the ferroelectric database of Smidt et al. and propose 6 candidates present in the Materials Project. Our results suggest new pathways to engineer small-band gap large-polarization ferroelectrics.
@@ I Acknowledgments
This work was supported by ANID FONDECYT Regular grant number 1220986. N.F.-C. acknowledges partial financial support from the project InTec, Code “FRO2395”, from the Ministry of Education of Chile. Powered@NLHPC: This research was supported by the supercomputing infrastructure of the NLHPC (CCSS210001).